The period of a 2 kg stone is rotated in a vertical circle on a 2 m cord with the period of 1 s.

 

Give the magnitude of the radial acceleration in m/s2.



1
Expert's answer
2021-04-13T06:42:14-0400

"a=\omega^2r,"

"\omega=\frac{2\pi}{T},"

"a=\frac{4\pi^2r}{T^2},"

"a=\frac{4\cdot 3.14^2\cdot 2}{1^2}=79~\frac{m}{s^2}."